OpenStudy (anonymous):

x^3-4x^2+8 use rational root theorem . I got the number two but I can't figure out the other two numbers

myininaya (myininaya):

are you only suppose to find the rational zeros?

myininaya (myininaya):

2| 1 -4 0 8 | 2 -4 -8 ----------------- 1 -2 -4| 0 you can find other 0's by solving \[x^2-2x-4=0\]

OpenStudy (anonymous):

yeah! oh duh! thanks so much!

OpenStudy (anonymous):

uaos;j whf except i don't know what im doing and now can't figure the rest out!

myininaya (myininaya):

use the quadratic formula

OpenStudy (anonymous):

ohhh . thanks!

OpenStudy (anonymous):

pardon? we can't apply quadratic formula given that the equation is in third degree.. it's a polynomial. . . . .

myininaya (myininaya):

\[x^2-2x-4=0\] is 2nd degree

myininaya (myininaya):

\[x=\frac{-(-2) \pm \sqrt{(-2)^2-4(1)(-4)}}{2(1)}\]

myininaya (myininaya):

\[x=\frac{2 \pm \sqrt{4+16}}{2}=\frac{2 \pm \sqrt{20}}{2}=\frac{2 \pm 2 \sqrt{5}}{2}=1\pm \sqrt{5}\]

myininaya (myininaya):

so we have the three solutions the one you found which is x=2 and the other or irrational solutions (the ones i found)

OpenStudy (anonymous):

in continuation to myinaniya. . . therefore; \[(x ^{2}-2x-4)(x-2)\] this are the factors derived from the original equation. to continue; to make a perfect square root \[(x ^{2}-2x-4)\] - to factor; \[x ^{2}-2x= 4\]\[x ^{2}-2x+1=4+1\]\[x ^{2}-2x+1=5\] \[(x-1)(x-1)=5\] the factors are: (x-2); (x-1)(x-1)=5 to solve: x-2=0 x-1=5 x-1=5 x=2 x=6 x=6 therefor roots are; (2,6,6)

myininaya (myininaya):

no you solve x^2-2x-4=0 wrong

myininaya (myininaya):

you can only do: if you have a*b=0, then either a=0 or b=0 or both=0 ------------------- \[(x-1)(x-1)=5\] \[(x-1)^2=5\] however you can solve this by using the definition of absolute value you need to take square root both sides to do this \[x-1=\pm \sqrt{5}\] now add 1 on both sides \[x=1 \pm \sqrt{5}\]

OpenStudy (anonymous):

thank you for the correction. . . .